OpenStudy (anonymous):

simplify. 6+2^2/52-2*6 SHOW ALL WORK

OpenStudy (anonymous):

is this like this (6+2^2)/(52-2*6 ) =(6+4)/(52-12) =10/40 =1/4 ans